# Traffic-R1: Reinforced LLMs Bring Human-Like Reasoning to Traffic Signal Control Systems

### v1

@@ 4.1.2. Offline Policy Optimization
Inspired by the work of (Liu et al., 2024), which effectively finetunes LLMs on mathematics and coding datasets containing only final answers or actions without reasoning text, our approach uses the same offline RL framework to finetune the LLM on our expert-collaborative TSC dataset to promote self-thinking in traffic control. The core process involves an LLM policy model interacting with an offline environment, guided by a rule-based reward derived from expert-provided answers. Given a textual prompt $x$ describing a specific traffic scenario, the LLM, parameterized by $\theta$, generates an output sequence $y$ autoregressively according to its policy $\pi_{\theta}(y\mid x)=\prod_{t=1}^{|y|}\pi_{\theta}(y_{t}\mid x,y_{<t})$.
@@
Reinforcing Reasoning via Policy Optimization. In the offline environment, the reward function $R$ evaluates the quality of output $y$. To optimize the policy parameters $\theta$, we employ Group Relative Policy Optimization (GRPO) (Liu et al., 2024) to ensure stable gradient updates. Let $\pi_{\theta}$ represent the policy and $\mathcal{B}={(x_{i},{a_{i}^{(j)}}_{j=1}^{k})}$ denote a batch of input prompts $x_{i}$, each paired with $k$ candidate completions $a_{i}^{(j)}$ sampled from the policy. The reward function $R$ assigns a score $r_{i}^{(j)}$ to each completion $a_{i}^{(j)}$. To address high variance in policy gradients, GRPO computes group-normalized advantages for each completion $a_{i}^{(j)}$ generated from the same input $x_{i}$, as shown in Equation 2. This approach centers the rewards within each group, mitigating the impact of absolute reward magnitudes:
(2) $A_{i}^{(j)}=r_{i}^{(j)}-\frac{1}{k}\sum_{l=1}^{k}r_{i}^{(l)},$
@@
$\mathcal{L}(\theta)$ $=\mathbb{E}_{(x_{i},a_{i}^{(j)}\sim\pi_{\theta_{\text{ref}}})}\Bigg{[}\min\left(\rho_{i}^{(j)}A_{i}^{(j)},\text{clip}(\rho_{i}^{(j)},1-\epsilon,1+\epsilon)A_{i}^{(j)}\right)$
(3) $\quad-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(\cdot|x)\parallel\pi_{\theta_{\text{ref}}}(\cdot|x)\right]\Bigg{]},$
@@
where $\rho_{i}^{(j)}=\frac{\pi_{\theta}(a_{i}^{(j)}|x_{i})}{\pi_{\theta_{\text{ref}}}(a_{i}^{(j)}|x_{i})}$ is the likelihood ratio between the current policy $\pi_{\theta}$ and the reference policy $\pi_{\theta_{\text{ref}}}$, and $\epsilon$ is the clipping threshold. The coefficient $\beta$ determines the strength of the Kullback-Leibler divergence penalty $\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\theta_{\text{ref}}}\right]$. In practice, the reference policy $\pi_{\theta_{\text{ref}}}$ is typically set to a snapshot of the previous policy, which stabilizes training to inspiring deep thinking instead of imitation by constraining policy updates.
4.2. Open-World Reinforcement Learning
@@
Although the human-informed offline RL has finetuned LLM to learn from human experts for stable performance on TSC tasks, the model’s capacity is limited to the expert knowledge extracted from the dataset. In this section, we propose an open-world online RL to inspire LLM explore multi-step and multi-intersection TSC networks by itself. This approach allows LLM to interact with the online dynamic simulated traffic environment and update its policy based on online reward for better performance.
@@
Online Traffic Environment Simulation To simulate the multi-intersection and multi-step dynamics of real-world traffic flow, we constructed a $\times 4$ 4 simulated road network with 300-meter roads between each intersection. The 16 positions in the network represent most typical road scenarios encountered at real-world intersections. Traffic flow within the network is randomly generated, allowing for up to 8,000 vehicles over the course of one hour. For efficiency in iterations, we utilize CityFlow (Zhang et al., 2019) as the simulator to model the traffic dynamics resulting from the actions of the LLM. The online multi-step rewards $R_{t}raj$ are quantified based on the cumulative average queue length and waiting time caused by a series of multi-step, multi-intersection actions. We use group advantage as final rewards to mitigate random variations in environment during online training through group mean comparison, ensuring stable optimization gradients.
Multi-step Policy Optimization Existing RL finetuning often relies on an offline single-turn setting, such as mathematical problem solving. However, our online RL approach requires LLMs to operate in interactive environments that evolve over multiple steps with stochastic feedback. Although (Wang et al., 2025b) proposed trajectory-level optimization for multi-step online scenarios by concatenating environment observations and LLM responses step by step with a trajectory reward, this solution is not suitable for traffic control tasks. The full trajectory for TSC observation and reasoning may become excessively long, resulting in high computation costs and sparse token-level attention for key tokens. Moreover, real-world TSC cannot be directly modeled as a perfect Markov Decision Process, as is done in digital games like (Wang et al., 2025b). Attempting to directly adapt the approach from (Wang et al., 2025b) to real-world TSC would likely yield results similar to those of traditional RL methods, which are characterized by poor generalization and limited explanatory power.
@@
To address this issue, we propose a specific multi-step policy optimization method for traffic scenarios. As traffic flow varies randomly over time in real-world systems, the connections between previous actions and the subsequent observation and action are not strictly continuous and exhibit chaotic dynamics. Our approach, termed Stepwise Trajectory Policy Optimization (STPO), involves directly assigning the trajectory reward to each individual step instead of entire trajectories, corresponding to a specific observation $o_{t}$ and response $a_{t}$ pair. This method decomposes the long trajectory into segments, thereby reducing computational overhead and enabling denser reward signals at the step level. To formalize STPO, we define the reward assignment and policy update process with the following equations. The step-wise reward $r_{t}$ for observation $o_{t}$ and action $a_{t}$ is derived from the trajectory reward $R_{t}raj$ by distributing it proportionally across steps:
(4) $r_{t}={R_{\text{traj}}(o_{1:T},a_{1:T})}\cdot{T}^{-1},$
where $R_{t}raj(o_{1:T},a_{1:T})$ is the total trajectory reward over the sequence of observations $o_{1:T}$ and actions $a_{1:T}$, and T is the number of steps. The policy $pi_{t}heta(a_{t}|o_{t})$ is updated using the advantage $A_{t}$ similar to Equation 2. The policy optimization objective is:
(5) $\mathcal{L}_{\text{STPO}}(\theta)=\mathbb{E}_{(o_{t},a_{t}\sim\pi_{\theta_{\text{ref}}})}\left[\log\pi_{\theta}(a_{t}|o_{t})A_{t}\right],$
@@
where the rest computations on Kullback- Leibler divergence penalty and advantage clipping are similar to orginal GRPO in Equation 4.1.2.
4.3. Asynchronous Communication Network
Existing research on TSC systems typically relies on a synchronized parallel workflow assumption for simplicity, where all TSC agents at intersections update their observations and actions simultaneously at each time step. In practice, this synchronized approach becomes a limitation when addressing multi-intersection coordination. In real-world settings, coordination operates asynchronously, with one agent sending a message, another receiving it along with an action and response, and the process continuing back and forth.As a result, implementations of multi-intersection coordination within existing synchronized parallel workflows are often inefficient, requiring extensive shared global observations for every intersection, or ineffective, leading to incompatible final decisions even with shared observations. While some reinforcement learning methods attempt to improve performance by incorporating global or neighborhood observations into intersection representations, their generalization and scalability remain problematic.
@@ 4.3. Asynchronous Communication Network
Figure 4. Introduction of Asynchronous Communication Network compared with conventional synchronized network
@@
To achieve natural multi-intersection communication and coordination, and to unlock the potential of the reinforced LLM TSC model, we design an asynchronous communication network for real-world multi-intersection TSC systems. As illustrated in Figure 4, we organize intersections into two groups based on the parity of their positions. Each conventional TSC step is divided into two inner steps: the first inner step activates models from group one, while the second inner step activates models from group two. This parity ensures that signal control across intersections operates in a structured spatial and temporal order. This arrangement allows models activated in the first inner step to generate messages for their neighboring intersections, which are then activated in the subsequent inner step to receive and process these messages. With this network, the language communication capabilities of LLMs can be fully utilized to enhance TSC systems.
Table 1. Statistics of traffic flow datasets.
@@
5. Experiments
Table 2. Zero-shot performance comparison on conventional traffic signal control tasks (the smaller the better). The best results are in bold and second-best results are underlined.
@@ Methods
  • RQ1: Can Traffic-R1 outperform other traffic signal control (TSC) models on public datasets and in zero-shot settings?
  • RQ2: How does Traffic-R1 perform in handling out-of-distribution incidents through human-like reasoning?
@@
5.1. Experimental Settings
5.1.1. Dataset
@@
Our experiments were primarily conducted on two public traffic flow datasets (Mei et al., 2024) to ensure fair comparison, as detailed in Table 1. The Jinan dataset comprises 12 intersections from Jinan, China, and includes three distinct traffic flow recordings representing different time periods. The Hangzhou dataset consists of 16 intersections from Hangzhou, China, with two recordings. For out-of-distribution scenarios, we collect traffic emergency incident recordings from traffic management departments and summarize them into 200 representative textual examples, such as passages running onto roads, vehicle accidents, and school times, along with the action records implemented by traffic managers as correct responses (presented in Appendix A.2). Besides, for emergency vehicle coordination scenarios, we modify the Hangzhou1 datasets by incorporating a 5% proportion of emergency vehicle flow to simulate real-world conditions.
5.1.2. Implementation Details
@@
We use CityFlow (Zhang et al., 2019), a widely used experimental simulator for traffic control, to test all models. For each dataset, we apply a standard action space consisting of four signal control phases: NTST (north-south through), ETWT (east-west through), ELWL (east-west left-turn), and NLSL (north-south left-turn). In the experimental environment, right-turn movements are permitted at all times, and each phase lasts for 15 seconds, aligning with typical requirements of real-world TSC systems (Zhang et al., 2022; Mei et al., 2024). Every phase action is followed by a three-second yellow phase and a two-second red phase to facilitate the transition to the next phase. All traffic flow datasets in experiments are simulated for one hour.
5.1.3. Baseline Methods
@@
We incorporate a range of baseline models from various research areas to ensure a comprehensive comparison. For traditional TSC methods, we include FixedTime (Koonce et al., 2008) and Maxpressure (Varaiya, 2013). For RL-based methods, we evaluate five effective approaches: MPLight (Chen et al., 2020), AttendLight (Oroojlooy et al., 2020), CoLight (Wei et al., 2019), Efficient-CoLight (Wu et al., 2021), and Advanced-CoLight (Zhang et al., 2022), along with the state-of-the-art LLM-based method, LLMLight (Lai et al., 2023). For zero-shot methods, we assess the performance of CoLLMLight (Yuan et al., 2025) and general LLM models, which include Llama 3.3 (70B), Qwen 2.5 (72B), GPT 3.5-turbo, GPT-4o, and DeepSeek-R1 (671B and distilled to 7B). All learning-based baselines are trained on the same 4x4 simulated road network and traffic flow dataset as Traffic-R1 during the open-world RL stage. Notably, for LLMLight and CoLLMLight, we also incorporate our expert-collaborative dataset into the training instructions to ensure a fair comparison
5.1.4. Evaluation Protocols
@@
For standard TSC tasks, we adopt the commonly used Average Travel Time (ATT) and Average Waiting Time (AWT) to evaluate the performance of models on conventional TSC scenarios. Lower values in ATT and AWT indicate better traffic efficiency brought by the models’ strategies.
5.2. Results on Conventional TSC Datasets (RQ1)
@@
We evaluate the performance of Traffic-R1 on conventional TSC tasks using public datasets that are widely adopted in TSC research. All learning-based methods are trained in the same simulated traffic environment as Traffic-R1 to fairly assess their zero-shot performance. The results, presented in Table 2, indicate that Traffic-R1, in a zero-shot configuration, outperforms all baselines by a significant margin, demonstrating strong cross-dataset generalization capability, which is advantageous for real-world deployment. Additionally, we observe poor zero-shot performance for RL-based methods, with results that are even worse than those of traditional methods. This underperformance underscores their unsuitability for real-world deployment, where there are no conditions for them to iterate and train. It is noteworthy that some advanced LLMs achieve impressive zero-shot results due to their deep reasoning abilities, such as DeepSeek-R1-671B; however, their performance significantly decreases when distilled into a smaller parameter size, such as DeepSeek-R1-Distill-7B. We also report the full-shot comparison results in Appendix A.4, where all RL-based models are trained on every test dataset following their original settings. In this setting, RL-based methods demonstrate strong advantages over traditional methods by continuously searching for better policies. Nevertheless, zero-shot Traffic-R1 still outperforms them through its internal human-like reasoning and traffic control policies.
5.3. OOD Incident Dealing Performance (RQ2)
Focusing solely on performance in ideal simulated environments is inadequate and unreliable for real-world deployment. In RQ2, we investigate whether Traffic-R1 can effectively handle OOD traffic scenarios through its human-like reasoning. Unlike conventional tasks, incident handling requires models to utilize internal logic, commonsense, and a deep understanding of TSC, presenting challenges for traditional RL-based models and even LLM-based methods.
We categorize OOD tasks into two types: 1. local intersection incidents, which occur at a single intersection and do not directly affect others (e.g., a traffic accident), and 2. network-wide incidents, which involve emergency vehicles that impact the entire network and require coordinated responses across multiple intersections (e.g., an ambulance navigating through the network). For local intersection incidents, we use Emergency Action Accuracy (EAA), defined as the ratio of correct actions taken by the model within the emergency incident dataset, to evaluate the model’s ability to respond accurately to incidents. For network-wide incidents, we adapt Average Emergency Travel Time (AETT) and Average Emergency Waiting Time (AEWT) for emergency vehicles from (Wang et al., 2024) to assess model effectiveness. We evaluate all models based on these two types of incidents, with results presented in Table 3. The findings highlight two key strengths of Traffic-R1:
  • Stable OOD scenario generalization. The proposed Traffic-R1 achieves stable performance across various OOD scenarios, surpassing most general LLMs, which are larger and slower to deploy. Specifically, Traffic-R1 outperforms LLMLight, a model instruction finetuned for TSC, by a significant margin of over 30% across all metrics. This demonstrates that Traffic-R1’s human-like reasoning effectively comprehend traffic knowledge and adapts to diverse scenarios, rather than repeat and imitation.
  • Lightweight and efficient. With only 3 billion parameters, Traffic-R1 achieves stable performance on these tasks, surpassing the capabilities of traditional baselines and matching advanced LLMs while using only 1% of their parameters and significantly lower deployment requirements. Notably, Traffic-R1 excels in handling network-wide incidents requiring multi-intersection coordination, demonstrating that our asynchronous communication network effectively enables the LLM-based TSC model to perform tacitly in complex, multi-intersection TSC systems.

### v2

@@ 4.1.2 Offline Policy Optimization
Inspired by the work of Liu et al. (2024), which effectively finetunes LLMs on mathematics and coding datasets containing only final answers or actions without reasoning text, our approach uses the same offline RL framework to finetune the LLM on our expert-collaborative TSC dataset to promote self-thinking in traffic control. The core process involves an LLM policy model interacting with an offline environment, guided by a rule-based reward derived from expert-provided answers. Given a textual prompt $x$ describing a specific traffic scenario, the LLM, parameterized by $\theta$, generates an output sequence $y$ autoregressively according to its policy $\pi_{\theta}(y\mid x)=\prod_{t=1}^{|y|}\pi_{\theta}(y_{t}\mid x,y_{<t})$.
@@
Reinforcing Reasoning via Policy Optimization. To optimize the policy $\pi_{\theta}$ with parameters $\theta$ in an offline environment, we employ Group Relative Policy Optimization (GRPO) Liu et al. (2024) to ensure stable gradient updates. Let $\pi_{\theta}$ represent the policy and $\{a_{i}^{(j)}\}_{j=1}^{k}$ denote each input prompt $x_{i}$ is paired with $k$ candidate completions sampled from the current policy. The reward function $R$ assigns a score $r_{i}^{(j)}$ to each completion $a_{i}^{(j)}$. To address high variance in policy gradients, GRPO computes group-normalized advantages for each completion $a_{i}^{(j)}$ generated from the same input $x_{i}$, as shown in Equation 1. This approach centers the rewards within each group, mitigating the impact of absolute reward magnitudes:
$A_{i}^{(j)}=r_{i}^{(j)}-\frac{1}{k}\sum_{l=1}^{k}r_{i}^{(l)},$ (1)
@@
$\begin{split}\mathcal{L}(\theta)=\mathbb{E}\Bigg[&\min\bigg(\rho_{i}^{(j)}A_{i}^{(j)},\text{clip}(\rho_{i}^{(j)},1-\epsilon,1+\epsilon)*\\ &A_{i}^{(j)}\bigg)-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(\cdot|x)\parallel\pi_{\theta_{\text{ref}}}(\cdot|x)\right]\Bigg],\end{split}$} (2)
@@
where $\rho_{i}^{(j)}=\frac{\pi_{\theta}(a_{i}^{(j)}|x_{i})}{\pi_{\theta_{\text{ref}}}(a_{i}^{(j)}|x_{i})}$ is the likelihood ratio between the current policy $\pi_{\theta}$ and the reference policy $\pi_{\theta_{\text{ref}}}$, and $\epsilon$ is the clipping threshold. The expectation $\mathbb{E}[\cdot]$ is computed over $(x_{i},a_{i}^{(j)})$ drawn from $\pi_{\theta_{\text{ref}}}$.The coefficient $\beta$ determines the strength of the Kullback-Leibler divergence penalty $\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\parallel\pi_{\theta_{\text{ref}}}\right]$. In practice, reference policy $\pi_{\theta_{\text{ref}}}$ is typically set to a snapshot of the previous policy, which stabilizes training to inspire deep thinking instead of imitation by constraining policy updates.
4.2 Open-World Reinforcement Learning
@@
Although the human-informed offline RL has finetuned LLM to learn from human experts for stable performance on TSC tasks, the model’s capacity is limited to the expert knowledge extracted from the dataset. In this section, we propose an open-world online RL to inspire LLM explore multi-step and multi-intersection TSC networks. This approach allows LLM to interact with the online dynamic simulated traffic environment and update its policy based on online reward for better performance.
@@
Online Traffic Simulation To simulate the multi-intersection and multi-step dynamics of real-world traffic flow, we constructed a $4\times 4$ simulated road network with 300-meter roads between each intersection. The 16 positions in the network represent most typical road scenarios encountered at real-world intersections. Traffic flow within the network is randomly generated, allowing for up to 8,000 vehicles over the course of one hour. For efficiency in iterations, we utilize CityFlow Zhang et al. (2019) as the simulator to model the traffic dynamics resulting from the actions of the LLM. The online multi-step rewards $R_{traj}$ are quantified based on the cumulative average queue length and waiting time caused by a series of multi-intersection actions. We use group advantage as final rewards to mitigate random variations in environment during online training through group mean comparison, ensuring stable optimization gradients.
Multi-step Policy Optimization Existing RL finetuning typically uses offline, single-turn settings (e.g., math). Traffic online RL requires LLMs to act in interactive, multi-step environments with stochastic feedback. Methods Wang et al. (2025b) that concatenate observations and model responses into a single trajectory with a trajectory-level reward are impractical for TSC: context lengths become very large, raising compute and memory costs and diluting token-level attention on critical observations and actions. Besides, traffic is also partially observable and not well modeled as a perfect Markov decision process, so directly porting trajectory-concatenation approaches from digital games will likely produce brittle policies that generalize poorly and offer limited interpretability, similar to traditional RL controllers.
@@
To address the chaotic dynamics in real-world traffic systems, where the link between actions and subsequent states is not strictly continuous, we propose Stepwise Trajectory Policy Optimization (STPO). This method assigns the total trajectory reward to each individual observation-action pair ($o_{t},a_{t}$), decomposing the trajectory to provide denser, step-level reward signals and reduce computational overhead. The step-wise reward $r_{t}$ is uniformly distributed from the total trajectory reward $R_{\text{traj}}$ over $T$ steps:
$r_{t}={R_{\text{traj}}(o_{1:T},a_{1:T})}\cdot{T}^{-1},$ (3)
where $R_{\text{traj}}$ is the reward over the full sequence of observations $o_{1:T}$ and actions $a_{1:T}$. The policy $\pi_{\theta}(a_{t}|o_{t})$ is then optimized with the objective:
$\mathcal{L}_{\text{STPO}}(\theta)=\mathbb{E}_{(o_{t},a_{t}\sim\pi_{\theta_{\text{ref}}})}\left[\log\pi_{\theta}(a_{t}|o_{t})A_{t}\right],$ (4)
@@
The advantage calculation ($A_{t}$), Kullback-Leibler divergence penalty, and advantage clipping are analogous to those in the original GRPO formulation (see Equation 2).
4.3 Asynchronous Communication Network
Most TSC research assumes a synchronized parallel workflow, where all intersection agents act simultaneously. This simplification hinders real-world multi-intersection coordination, which is inherently asynchronous. Consequently, implementations in synchronized frameworks are often inefficient, requiring extensive shared global data, or ineffective, leading to incompatible decisions. While some reinforcement learning methods incorporate neighborhood observations, they typically suffer from poor generalization and scalability.
@@ 4.3 Asynchronous Communication Network
Figure 4: Asynchronous communication network compared with conventional synchronized network
@@
To enable natural communication and unlock the potential of our LLM-based model, we propose an asynchronous communication network. As illustrated in Figure 4, we partition intersections into two groups based on positional parity. Each TSC step is split into two inner steps: the first activates group one, and the second activates group two. This structure allows models activated in the first inner step to generate messages for their neighbors, which are then received and processed in the subsequent inner step. This framework fully utilizes the language capabilities of LLMs to enhance multi-intersection coordination. The pseudocode for this network is provided in Algorithm 1.
5 Experiments
Table 1: Statistics of traffic flow datasets.
@@ 5 Experiments
[…]
@@ Methods
  • RQ1: Can Traffic-R1 outperform other TSC expert models and LLMs on public datasets and in zero-shot settings?
  • RQ2: How does Traffic-R1 perform in handling OOD incidents through its reasoning?
@@
5.1 Experimental Settings
5.1.1 Dataset
@@
Our experiments were primarily conducted on two public traffic flow datasets Mei et al. (2024) to ensure fair comparison, as detailed in Table 1. For out-of-distribution scenarios, we collect traffic emergency incident recordings from traffic management departments and summarize them into 200 representative textual examples, such as passages running onto roads, vehicle accidents, and school times, along with the action records implemented by traffic managers as correct responses (presented in Appendix A.3). Besides, for , we modify the Hangzhou1 datasets by incorporating a 5% proportion of emergency vehicle flow to simulate emergency vehicle coordination scenarios.
5.1.2 Implementation Details
@@
Traffic-R1 was trained based on Qwen2.5-3B-base on a device with 4 H100 GPUs and deployed for inference on a single Tesla T40 GPU, utilizing the Verl framework Sheng et al. (2024). We evaluated all models using the CityFlow traffic simulator Zhang et al. (2019). The experimental setup features a standard four-phase action space: north-south through (NTST), east-west through (ETWT), east-west left-turn (ELWL), and north-south left-turn (NLSL). Each green phase lasts 15 seconds, followed by a 3-second yellow and a 2-second red transition period, consistent with real-world TSC systems Zhang et al. (2022); Mei et al. (2024). Right turns are permitted at all times. All traffic flow datasets are simulated for one-hour periods.
5.1.3 Baseline Methods
@@
We incorporate a range of baseline models from various research areas to ensure a comprehensive comparison. For traditional TSC methods, we include FixedTime Koonce et al. (2008) and Maxpressure Varaiya (2013). For RL-based methods, we evaluate five effective approaches: MPLight Chen et al. (2020), AttendLight Oroojlooy et al. (2020), CoLight Wei et al. (2019), Efficient-CoLight Wu et al. (2021), and Advanced-CoLight Zhang et al. (2022), along with the state-of-the-art LLM-based method, LLMLight Lai et al. (2023). For zero-shot methods, we assess the performance of CoLLMLight Yuan et al. (2025) and general LLM models, which include Llama 3.3 (70B), Qwen 2.5 (72B), GPT 3.5-turbo, GPT-4o, and DeepSeek-R1 (671B and distilled to 7B). All learning-based baselines are trained on the same 4x4 simulated road network and traffic flow dataset as Traffic-R1 during the open-world RL stage. Notably, for LLMLight and CoLLMLight, we also incorporate our expert-collaborative dataset into the training instructions to ensure a fair comparison
5.1.4 Evaluation Protocols
@@
We adopt the commonly used Average Travel Time (ATT) and Average Waiting Time (AWT) to evaluate the performance of models. Lower values in ATT and AWT indicate better traffic efficiency.
5.2 Conventional TSC Evaluation (RQ1)
@@
We evaluate the performance of Traffic-R1 on conventional TSC tasks using public datasets that are widely adopted in TSC research. All learning-based methods are trained in the same simulated traffic environment as Traffic-R1 to fairly assessment. As shown in Table 2, our model significantly outperforms all baselines, demonstrating strong generalization for real-world deployment. RL-based methods perform poorly in zero-shot scenarios, even lagging behind traditional methods, which questions their real-world applicability without iterative training. While large LLMs like DeepSeek-R1-671B achieve impressive zero-shot results, their performance degrades when distilled into smaller models. For completeness, Appendix A.4 presents full-shot results, where RL models are trained on each test dataset. Although RL methods then surpass traditional ones, our zero-shot model still outperforms them, highlighting its superior internal reasoning and control policies.
5.3 OOD Incident Dealing (RQ2)
Evaluation only in ideal simulations is insufficient for deployment. RQ2 asks whether Traffic-R1 can handle out-of-distribution (OOD) traffic incidents via human-like reasoning. Incident handling requires internal logic, commonsense, and domain knowledge, which challenge both traditional RL and standard LLM approaches.